### v2

# Improving Model Evaluation using SMART Filtering of Benchmark Datasets

Abstract
One of the most challenging problems facing NLP today is evaluation. Some of the most pressing issues pertain to benchmark saturation, data contamination, and diversity in the quality of test examples. To address these concerns, we propose Selection Methodology for Accurate, Reduced, and Targeted (SMART) filtering, a novel approach to select a high-quality subset of examples from existing benchmark datasets by systematically removing less informative and less challenging examples. Our approach applies three filtering criteria, removing (i) easy examples, (ii) data-contaminated examples, and (iii) examples that are similar to each other based on distance in an embedding space. We demonstrate the effectiveness of SMART filtering on three multiple choice QA datasets, where our methodology increases efficiency by reducing dataset size by 48% on average, while increasing Pearson correlation with rankings from ChatBot Arena, a more open-ended human evaluation setting. Our method enables us to be more efficient, whether using SMART filtering to make new benchmarks more challenging or to revitalize older datasets, while still preserving the relative model rankings.
Improving Model Evaluation using SMART Filtering of Benchmark Datasets
Vipul Gupta1,2,121,2$\dagger$1 , 2 , †  Candace Ross1111  David Pantoja3333  Rebecca J. Passonneau2222 Megan Ung1111Adina Williams1111 $1$ FAIR, Meta AI  $2$ Pennsylvania State University  $3$ University of California, Berkeley vkg5164@psu.edu, {ccross, meganu, adinawilliams}@meta.com
$\dagger$ This work was done during an internship at FAIR, Meta.
1 Introduction
Figure 1: This figure illustrates our methodology. We remove easy, data contaminated, and/or similar examples from datasets to find a high-quality subset. The examples selected are from MMLU dataset.
Dataset Filtering Steps (in %) SMART Filtering
Name # of examples Easy Data Contaminated Similar Prefiltering % Filtered # examples
ARC 3530 64.41 3.45 12.57 0.22 68.92 1097
MMLU 14042 35.01 4.37 3.31 7.61 43.02 8000
CommonsenseQA 1221 27.60 0.00 8.90 0.16 34.30 802
Table 1: SMART filtering of MMLU, CommonsenseQA, and ARC results in leaner, more informative evaluations.
With the recent rise of interest in natural language generation and language modeling as a general purpose pretraining task, model benchmarking—evaluating several models on a standard test set to derive a strict ranking—has become much less straightforward. This is in part due to the increasing speed of evaluation dataset saturation (Kiela et al., 2021; Vania et al., 2021; Ott et al., 2022), whereby current state-of-the-art models are able to match human-level performance within the margin of error Laskar et al. (2023); Keegan (2024). This signals the need for new, potentially more challenging benchmarks; however, generating high quality human-annotated datasets is difficult Rein et al. (2023). It is time-consuming to devise new evaluation datasets, which can require training and coordinating human annotators, cleaning data, and assessing test validity Sun et al. (2021); Herde et al. (2021). Moreover, the increasing speed of benchmark saturation leads to a difficult value proposition for benchmark creators who have only so much time to make important research contributions.
As if the speed of progress in NLP benchmark saturation weren’t enough of a challenge, the increasingly wide availability of text-generation systems has made the classical approach to collecting high quality, human-baselined evaluation data—crowd-sourcing—much less tenable. Since detecting model-generated text is an open area of research (Jawahar et al., 2020; Dugan et al., 2023), benchmark developers can no longer trust that a new, purportedly human-created evaluation dataset was indeed created by humans, and not humans using automatic tools as assistants. This state of affairs raises an important question about past evaluation datasets: if they weren’t saturated, could they still help us rank models based on task performance? Can we recapture some of the utility in saturated evaluation datasets that appear to be saturated?
One common approach to dataset saturation is filtering Bras et al. (2020). Just as renovating an older home is generally more efficient than building a new home from scratch, dataset filtering can rejuvenate a high-quality older benchmark, without the need to rely on novel rounds of costly human data collection. Along these lines, we propose Selection Methodology for Accurate, Reduced and Targeted (SMART) filtering, a novel approach for finding a high-quality subset of evaluation benchmarks that doesn’t need human intervention, illustrated in Figure 1. We achieve this by algorithmically removing the low-quality examples from existing evaluation datasets, preserving the high-quality examples that are most informative and reliable as our basis for benchmarking. Our methodology applies three filtering criteria: 1) removing easy examples, 2) removing data-contaminated examples that are highly likely to have been leaked into the training datasets, and 3) removing similar examples.
All SMART filtering steps have their conceptual origin in dataset filtering, which can reduce the data size and hence compute costs (Treviso et al., 2023; Mishra and Sachdeva, 2020). In particular, we utilize filtering methods used for training-time deduplication (Lee et al., 2022; Abbas et al., 2023; Swayamdipta et al., 2020), a technique often used during pretraining to increase efficiency. We contend that pretraining deduplication approaches will also be effective for benchmarking filtering. They can increase headroom on standard, already vetted, human-written datasets that have been saturated, and re-enable meaningful model rankings.
Dataset filtering, if correctly applied (c.f., Phang et al. 2022), can retain or even improve benchmark dataset quality (Treviso et al., 2023), meaning that SMART filtering might be able to contribute to addressing reliability issues arising from current day benchmarking practices Gupta et al. (2024); Alzahrani et al. (2024); Sinha et al. (2021). Each of the steps in SMART filtering aims to improve usefulness and reliability by removing examples that are not informative for differentiating models.
SMART filtering could also be applied to new datasets at the earlier stages of the benchmark lifecycle before they become standards. When used before new benchmarks premiere, SMART filtering can ensure that the final version is as challenging, high quality, and efficient as possible.
We test SMART filtering on the classical multiple choice QA setting. We characterize the impact of SMART filtering using three datasets—ARC Clark et al. (2018), MMLU Hendrycks et al. (2021a) and CommonsenseQA Talmor et al. (2019). In Tables 1 and 2, we show that our methodology increases efficiency without hits to leaderboard utility. For example, on the ARC dataset, we shrink the dataset by 68.9%, increasing computational efficiency drastically, while observing practically no effect on the relative ranking of models. We also compare model rankings of a dataset with open-ended human evaluated ChatBot Arena Chiang et al. (2024) rankings, and found that datasets using SMART filtering have a stronger correlation with ChatBot Arena indicating better correlation with human preferences.
Qwen2-72B-it
Llama-3.1-70B-it
Llama-3-70B-it
Gemma-2-27b-it
Phi-3.5-MoE-it
Mixtral-8x22B-it
Dbrx-it
Yi-34B
Llama-3-8B-It
Qwen-7B-Chat
Gemma-7b-it
ARC 93.9 (1) 93.4 (2) 93.3 (3) 92.5 (4) 92.0 (5) 91.6 (6) 90.5 (7) 90.2 (8) 88.3 (9) 76.4 (10) 73.8 (11)
ARC-SMART 83.0 (1) 81.9 (2) 81.9 (2) 78.8 (3) 78.5 (4) 76.2 (5) 73.2 (6) 72.4 (7) 72.1 (8) 51.8 (10) 53.1 (9)
MMLU 84.1 (1) 82.3 (2) 80.3 (3) 76.2 (6) 78.7 (4) 77.9 (5) 73.2 (8) 75.8 (7) 66.4 (9) 56.3 (10) 51.7 (11)
MMLU-SMART 74.3 (1) 71.4 (2) 69.2 (3) 63.9 (6) 67.0 (4) 65.3 (5) 60.0 (7) 62.4 (8) 50.5 (9) 41.5 (10) 38.9 (11)
CommonsenseQA 88.5 (1) 81.5 (3) 83.3 (2) 79.4 (6) 80.9 (4) 75.9 (9) 78.6 (7) 79.7 (5) 76.5 (8) 65.2 (11) 68.8 (10)
CommonsenseQA-SMART 84.5 (1) 74.1 (3) 77.1 (2) 71.9 (5) 73.9 (4) 67.2 (9) 70.4 (7) 71.8 (6) 68.0 (8) 55.7 (11) 59.4 (10)
Table 2: Model performance on ARC, MMLU, CommonsenseQA, and their SMART-filtered counterparts. Accuracy scores are accompanied by rankings in parentheses. Cells highlighted in blue indicate models whose ranking shifted after applying SMART filtering. The notation ‘it’ refers to instruction-tuned versions of the model.
2 Related Works
Recently, many benchmark datasets such as MMLU Hendrycks et al. (2021a), GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021b) and GPQA Rein et al. (2023) have been proposed to measure the capabilities of language models (LMs). However, recent works have highlighted issues that call into question the reliability of these datasets. Despite extensive efforts in dataset creation, annotation errors still persist Gema et al. (2024); Wang et al. (2024b). Recently, some works have shown that the accuracy of models on multiple-choice question datasets can change significantly by simply altering the order of answer options Pezeshkpour and Hruschka (2024); Gupta et al. (2024); Sugawara et al. (2020); Zong et al. (2024). One proposal suggests modifying the self-attention matrix to prevent answer options from paying attention to each other by masking their scores to zero (McIlroy-Young et al., 2024), but nonetheless, benchmark validity and gameability remains an open area of ongoing research.
Further complicating evaluation, some models have a prior bias towards a particular option id (e.g. ‘A’) Zheng et al. (2024); Wei et al. (2024); Zheng et al. (2023); Li and Gao (2024); Reif and Schwartz (2024); Ross et al. (2024). Additionally, models can perform well above random chance even when question text is removed and only answer options are given, potentially highlighting possible data contamination issues Balepur et al. (2024); Shah et al. (2020). Recent works have shown that replacing the correct option with “None of the above” leads to a drastic decline in performance across all models (Wang et al., 2024a; Xu et al., 2024). This calls into question the reliability of datasets and leaderboard ranking as the accuracy of a model may not be representative of its capabilities Röttger et al. (2024); Raj et al. (2023). Under the assumption that leaked, gameable, spuriously correlated, and/or biased examples are often easier for models to do well on, we aim to help address such issues by filtering “easier” examples from benchmarks to increase reliability.
Other works have explored related approaches to improve evaluation Vivek et al. (2024); Bras et al. (2020); Gupta et al. (2023). Varshney et al. (2022) tested models only on difficult examples of a dataset, showing that evaluating on just 5% examples can achieve a very high correlation with the full dataset. However, their approach requires retraining the models on various subsets of training data and uses humans to verify the filtered dataset. In the current NLP paradigm, access to training data is often unavailable for most models Longpre et al. (2024) and even when accessible, it is very expensive to retrain models on those datasets. Contrasting with Varshney et al. (2022), our approach emphasizes computational efficiency and identifies a high-quality subset that correlates well with the original dataset, without requiring access to training data or expensive human verification.
3 SMART Filtering Methodology
Our methodology has three main filtering steps, each of which is applied independently; the order of filtering steps does not impact the final subset.
All three steps in SMART filtering are aimed at deduplicating and making NLP leaderboarding more efficient. We contend that examples in the dataset that all tested models get correct with very high confidence are not useful for establishing model ranking (although they might be useful for data exploration or other things). Therefore, including such “easy” examples merely increases the computation cost without providing ability to distinguish between models in leaderboard settings Varshney et al. (2022). Relatedly, examples that are present in pretraining datasets should not be used for model testing (Elangovan et al., 2021). Leaked examples can give some models unwanted and difficult-to-interpret advantages when they are used for a leaderboard (Jiang et al., 2024; Ravaut et al., 2024). Finally, examples that are overly similar to one another effectively contribute less information about model performance than examples that meaningfully differ. Example overlap can also contribute towards idiosyncratically favoring particular kinds of information and potentially even double counting. In what follows, we will describe each step in the SMART filtering pipeline, to enable others to apply our methodology to benchmarks from any NLP task.
3.1 Pre-filtering
Before applying the SMART filtering steps, we apply two pre-filtering criteria. First, we eliminated exact match duplicate examples, keeping only one instance of each unique example. Copied examples in a dataset lead to overestimation of model performance Matatov et al. (2022), and thus they should be removed. We leave this step as a “pre-filtering” step, as most evaluation datasets have already been filtered for exact-match copies by their creators, usually making this step unnecessary. We provide more details in Appendix A.3.
Our second prefiltering criteria is the removal of anomalous example subsets, which we define as examples that are significantly different in form from the rest of the dataset, as this makes them incommensurable and hard to interpret. These subsets are often low quality and decrease the reliability of the dataset. For instance, the moral scenarios subset in MMLU, a subset of the ETHICS dataset (Hendrycks et al., 2020), is notably different from other categories. Moreover, automating machine morality has been argued to have deep conceptual issues (Talat et al., 2022), raising questions about whether such a task should be included in a general purpose benchmark. We removed such subsets.
3.2 Filtering Easy Examples
In this step, we focus on identifying and removing the easy examples from the dataset Gururangan et al. (2018); Rodriguez et al. (2021a). We define easy examples based on the agreement between top-performing open-source models from Open LLM leaderboard (HuggingFace, [2024]). Specifically, we consider an example to be easy if there is unanimous agreement among all top-performing models, with each model answering the example correctly with high confidence (greater than $0.8$). This approach finds examples that are consistently easy across different model architectures and models trained on different datasets. By removing these examples, we also reduce the computational cost of running inference on the resulting dataset.
The rationale behind removing these examples is that such agreement between various models indicates that they do not help in providing meaningful insights into relative model capabilities as all models answer them correctly. However, to ensure future models don’t lose the ability to answer these easy examples, we elected to retain a random 10% of the easy examples in the SMART filtering (and associated them with metadata to enable future analysis). This helps us to adhere more closely to the distribution of the original benchmarks, while still achieving significant computational efficiency.
3.3 Filtering Data Contaminated Examples
In this step, we identify and remove examples that are likely to be data contaminated, i.e. present in training data of models Deng et al. (2023); Elazar et al. (2024); Magar and Schwartz (2022). This is important because evaluating models on examples they have already seen during training can lead to inflated numbers, giving a false sense of higher model performance. Detecting data contamination is inherently challenging as also highlighted in recent studies Duan et al. (2024); Singh et al. (2024), and many promising approaches require access to training data Jiang et al. (2024). To identify data contaminated examples, we follow an approach inspired by Balepur et al. (2024) and Balepur and Rudinger (2024) as shown in Figure 2. In this approach, we do not need access to training data. Specifically, we modify the prompt to the model by removing the question text and presenting only the answer choices. This approach challenges the model to select the correct answer in an artificial setting without the context of the question itself.
Question: What is second most common element in solar system?
Answers: (A) Iron (B) Hydrogen (C) Methane (D) Helium
Question: What is second most common element in solar system?
Answers: (A) Iron (B) Hydrogen (C) Methane (D) Helium
Figure 2: We remove question context and give answer-only prompt. If all models still predict the correct answer with high probability ($>0.8$), then we categorize that example as data contaminated.
In this step, we again take agreement between top-performing models. We employ a stringent criterion: an example is deemed contaminated only if all top n open-source models answer it correctly without any context and with high confidence (greater than $0.8$). This conservative approach prioritizes precision over recall to ensure that the examples removed are truly contaminated. We recognize that this strict criterion may underestimate the number of contaminated examples. As the research in detecting data contamination improves, we plan to incorporate more sophisticated techniques that can detect contamination more effectively while balancing precision and recall.
3.4 Filtering Similar Examples
This filtering step identifies highly similar examples within the dataset to prevent redundant evaluations and avoid potential bias towards models that perform well on these particular closely related examples. We use an embedding-based approach, representing each example (question and answer choices) in high dimensional space. We experimented with two prominent embedding methods: SentenceBert Reimers and Gurevych (2019), which encodes each input into a $384$-dimensional vector, and LLM2Vec BehnamGhader et al. (2024), which modifies decoder-only LMs for sentence representation. Both approaches use bidirectional attention to capture rich contextualized representation. However, due to widespread usage in various applications Su et al. (2023); Khattab and Zaharia (2020), and lower computational requirements, we prioritized SentenceBert in SMART filtering.
Figure 3: Cosine distances between SentenceBert embeddings for MMLU examples. The black vertical line is the threshold for identifying similar example pairs.
We compute the cosine distance between each pair of example embeddings. Cosine distance is defined as the cosine similarity subtracted from $1$ and ranges from 00 to $2$, with 00 signifying maximum similarity between two embeddings. To determine an appropriate threshold, $\delta$, for identifying similar pairs, we analyze the distribution of cosine distances as shown in Figure 3. We employ kernel density estimation to locate the first local maximum of this distribution, capturing the point at which the similarity begins to decrease significantly. This data-driven approach to selecting $\delta$ allows us to adapt to the specific characteristics of each dataset. Pairs with distance lower than $\delta$ (indicated by the black vertical dotted line in Figure 3) are considered similar and grouped into clusters. To maintain dataset distribution, we randomly remove half of the examples from each identified cluster. We provide more details on threshold selection in Appendix A.1.
Figure 4: Heatmap illustrating the degree of agreement between model predictions for ARC, MMLU and Commonsense QA datasets, as well as their SMART filtering variants. Notably, SMART filtering leads to a decrease in inter-model agreement, indicating that our approach helps in better differentiating the capabilities of the models.
4 Results: Efficiency & Informativeness
In this section, we present the results of applying SMART filtering on popular datasets, focusing on three well-established multiple-choice question answering datasets: ARC Clark et al. (2018), MMLU Hendrycks et al. (2021a) and CommonsenseQA Talmor et al. (2019). While we show results on MCQA datasets, our methodology is generic and can be applied to various dataset types. Given that all other classification tasks can be, and often are, recast into MCQA format, it strikes us as likely that SMART-filtering would work just as well for classical NLU tasks (like NLI, coreference resolution, sentiment analysis) without needing novel strategies. Although these tasks have conceptual differences, with respect to the format, these variations boil down to having different numbers of answer options. Going into specifics of why we selected these datasets, we selected MMLU and ARC datasets because of their widespread adoption in benchmarking by state-of-the-art models Meta AI (2024); Google (2024) and their coverage of diverse topics and difficulties. These two datasets are 4-choice MCQA, so to broaden the scope, we included CommonsenseQA, which offers 5-choice MCQA. We think our inclusion of different numbers of MCQA answer options between 4 and 5 gives some indication that the methodology will be adaptable to smaller numbers of answer options (2 or 3) too. We release our code and provide SMART filtering version of each dataset111 Link to our codebase: https://github.com/facebookresearch/ResponsibleNLP/tree/main/SMART-Filtering..
To identify easy and data contaminated examples, we evaluate 7 open source models from the top of the Open LLM leaderboard (HuggingFace, [2024]) on each QA dataset: Llama 3.1-70B (Meta AI, 2024), Phi-3-14B (Abdin et al., 2024), Yi-34B ($01$AI et al., 2024), Qwen2-72B (Yang et al., 2024), Mixtral-8x22B (Mistral AI, 2023), Gemma2-27B (Google, 2024), and DBRX-132B (Databricks, 2024). We use instruction tuned (‘instruct’) versions when available, as they should have better question-answering capabilities.
Moreover, as these models came from different organizations, they were likely trained on different datasets. Ideally, this minimizes the risk of bias towards any particular model when identifying easy and contaminated examples (see Section 5.1 for more discussion on this). We used open-source models, as we aim to develop a methodology that is accessible to any AI practitioner with sufficient computation resources, thereby avoiding the substantial closed-source API costs.
The effectiveness of SMART filtering is shown in Table 1 as the percentage of examples filtered for ARC, [^2] Because we excluded the 37 questions with 5 answer options instead of 4, ARC dataset size is $3530$, not $3567$. MMLU, and CommonsenseQA. [^3] We used validation set as test set answers are not provided. The change in model performance after SMART filtering is shown in Table 2. Notably, the relative ranking of models remains almost unchanged across both the original and filtered datasets. However, our methodology significantly reduces the dataset sizes, achieving up to a 68.9% reduction for ARC, which directly correlates to a commensurate reduction in computational costs for model evaluation.
The impact of SMART filtering varies significantly across datasets. In ARC, 64.4% of the examples were classified to be ‘easy’, compared to only 27.6% in CommonsenseQA. In MMLU, a significant 4.37% of the dataset was found to be data contaminated, while CommonsenseQA had none. Additionally, 12.57% of ARC and 4.56% of MMLU examples were identified as similar. These findings highlight that different datasets might suffer from different types of low-quality examples. For MMLU, a widely accepted dataset for measuring LLM performance, we were able to filter out 43% of the dataset while maintaining almost the same model ranking. This suggests that MMLU might not be saturated yet, and there is still room for further hill-climbing for current models.
While relative model ranking is maintained, SMART filtering leads to a notable decrease in inter-model agreement across all datasets, as measured by the percentage of samples yielding identical predictions among models (Figure 4). This reduction in model agreement is beneficial, as it enables better differentiation among models based on their capabilities, and provides more headroom for improvement by decreasing their correlation.
In the following subsections, we present category-specific results for select datasets for detailed analysis, measure the correlation in model rankings before and after SMART filtering, and compare model accuracies with ChatBot Arena scores Chiang et al. (2024).
Dataset Pearson’s Corr. Kendall’s $\tau$
Orig. SMART
ARC 0.783 0.845 0.951
MMLU 0.764 0.776 0.978
CommonsenseQA 0.666 0.660 0.965
Table 3: Pearson correlation is high between Elo scores on ChatBot Arena and model accuracies. Kendall’s Tau correlation of model ranking on original and SMART filtering subsets is also high.
4.1 Category-Wise Results
Many datasets have multiple sub-categories. For instance, MMLU comprises 57 categories such as ‘high school mathematics’, while ARC is divided into ARC-Easy and ARC-Challenge.
Our analysis reveals significant variation in the percentage of examples removed by SMART filtering across different categories in MMLU. Some categories appear to be more challenging than others. Over 60% of examples were removed in 9 categories. Notably, ‘high school government and politics’, ‘high school psychology’, ‘marketing’, and ‘sociology’ saw a reduction in size by 73%, 67%, 63%, and 62% respectively. A qualitative look into examples of these categories indicates they contain questions with well-known answers, making them easier for models and more susceptible to be present during training and possible data contamination, while less than 20% were removed in $13$ categories. ‘Abstract algebra’, ‘global facts’, ‘high school physics’, and ‘professional law’ have only 4%, 4%, 5%, and 12.6% of examples removed, respectively. These subjects require complex reasoning and problem-solving, posing greater challenges for models. A detailed breakdown is shown in Appendix A.4.
State-of-the-art models often report accuracy on the ARC-Challenge category of ARC, as it is presumed to be more difficult Meta AI (2024); Google (2024). Surprisingly, our analysis shows no significant difference between ARC-Easy and ARC-Challenge after SMART filtering. Both categories were significantly reduced, with 73% of ARC-Easy and 60% of ARC-Challenge examples removed. A high proportion, 55.1% of ARC-Challenge were deemed easy by all models.
4.2 Correlation in Model Rankings
To quantitatively assess the consistency of model rankings between the original datasets and their filtered versions, we employed Kendall’s Tau correlation coefficient Kendall (1938). This metric ranges from $1$ (perfect positive correlation) to $-1$ (perfect negative correlation).
Our analysis of 29 models reveals high correlations between the ranking before and after SMART filtering. Table 3 shows the Kendall’s Tau correlation coefficients for ARC, MMLU and CommonsenseQA were $0.951$, $0.978$, and $0.968$, respectively. The main purpose of many evaluation datasets, particularly those used in leaderboards, is to compare relative model performance. Our approach shows similar model rankings while significantly reducing dataset size (as high as $68.9\%$ for ARC) and increasing headroom.
4.3 Correlation between ChatBot Arena scores and Model Accuracy
ChatBot Arena Chiang et al. (2024) is considered one of the most trusted current sources for assessing model performance (Saxon et al., 2024). They use human preference to assign an Elo score Elo (1967) for models on their leaderboard Chiang et al. ([2024]). Some researchers consider these scores a more accurate reflection of real-world model usage than traditional benchmarks (Thompson et al., 2020; Raji et al., 2021; Ott et al., 2022; Saxon et al., 2024). Developing high-correlation proxies for ChatBot Arena is valuable, as it could reduce the need for costly human preference collection Ni et al. (2024).
We examine the relationship between model accuracies and ChatBot Arena Elo scores to analyze how well SMART filtered datasets reflect human preference metrics. Using Pearson correlation Freedman et al. (2007), we measure the linear relationship between accuracies and Elo scores for $29$ tested models. We use Pearson correlation, because it accounts for differences in Elo scores and not just relative model rankings.
Dataset Name Original SMART Random IRE
ARC 0.783 0.845 0.774 0.784
MMLU 0.764 0.776 0.767 0.766
CommonsenseQA 0.666 0.660 0.659 0.658
Table 4: Comparison of performance metrics across different methods. Random signifies random baseline as mentioned in Section 4.4.
Correlations are high across the board between ChatBot Arena Elo scores and the original versions of ARC ($r=0.783$) and MMLU ($r=0.764$); see Table 3. In fact, after SMART filtering the correlations increase, suggesting that filtered ARC ($r=0.845$) and filtered MMLU ($r=0.776$) are even better proxies than the original datasets for human preference rankings. This highlights the effectiveness of SMART filtering in both computational efficiency and alignment to human preferences.
4.4 Comparison with Other Approaches
In this section we compare SMART filtering with other similar approaches and a random baseline. Rodriguez et al. (2021b) introduce a method that filters examples based on difficulty and discriminability, aiming to make leaderboards more informative. For this baseline, we used 25th percentile thresholds for difficulty and discriminability criteria to establish this baseline. We call their method IRE. We also provide analysis with a random baseline where we created a random baseline by reducing each dataset to the same size as our SMART-filtered datasets (reducing dataset size randomly to 69%, 43% and 34% for ARC, MMLU and CommonsenseQA).
Dataset Name % of Examples Filtered
SMART IRE
ARC 68.9 28.1
MMLU 43.0 31.1
CommonsenseQA 34.3 27.8
Table 5: Percentage of examples filtered by SMART vs. IRE for each dataset.
As seen in Table 4, SMART Filtering consistently achieves a higher correlation with human preferences than both the random baseline and IRE. Additionally, SMART Filtering consistently filters out a larger portion of the datasets compared to IRE, leading to greater efficiency in evaluation (c.f. Table 5).
5 Robustness of our Approach
In this section, we show that SMART filtering is robust to altering settings, such as (i) the number of models used for identifying easy and data-contaminated examples and (ii) choice of embedding model used to find similar examples.
# of Models % of Examples Filtered
ARC MMLU CommonsenseQA
4 $72.9\pm 3.1$ $49.6\pm 3.0$ $41.2\pm 4.3$
5 $72.4\pm 1.1$ $46.1\pm 1.6$ $36.4\pm 1.4$
6 $69.7\pm 1.1$ $44.8\pm 1.1$ $34.6\pm 1.1$
Table 6: Effect of varying number of models used for easy and data contamination filtering.
5.1 Ablation: Number of Models for Easy and Data-Contaminated Examples
In the SMART filtering methodology, we used 7 models for identifying easy and data-contaminated examples. Our selection was based on selecting the top 7 opensource models from different organizations. To assess the robustness of our approach, we conducted an ablation study varying the number of models used. We randomly selected subsets of 4, 5, and 6 models from the original set of 7. Specifically, we run our methodology on 10 random subsets of 4 models from a total of 35 subsets ($\binom{7}{4}=35$), 10 random subsets of 5 models from a total of 21 subsets ($\binom{7}{5}=21$) and all 7 subsets of 6 models ($\binom{7}{6}=7$). The results, presented in Table 6, show that the percentage of examples filtered remains relatively stable across different model combinations for all three datasets. This shows that our methodology is largely insensitive to the number of models used, and can be effectively applied using any n top performing open-source models, again opening up possibilities for efficiency gains.
5.2 Embeddings for Similar Examples
Embedding Pair Percentage Overlap
SBert & Llama-3-8B 89.3
SBert & Mistral-7B 88.1
Table 7: Percentage of examples with extremely high semantic similarity of SentenceBert with LLM2Vec embedding, based on a distance in the embedding space smaller than a $\delta$ (see Section 3.4 for more on $\delta$).
We use SentenceBert embeddings to capture the meaning of each example Reimers and Gurevych (2019). The choice was motivated by its bi-directional attention that helps capture rich representations. Recent efforts have explored transforming decoder-only LMs into bi-directional text encoders, such as LLM2Vec BehnamGhader et al. (2024). To validate the robustness of this approach regardless of embedding model choice, we compare the model we used, SentenceBERT, with another embedding method, LLM2Vec.
Our results in Table 7 show a high degree of similarity between the two embedding methods, with an average overlap of 88.7% between SentenceBERT and different LLM models. This suggests comparable accuracy in identifying similar examples for both SentenceBert and LLM2Vec, aligning with findings in Freestone and Santu (2024) that LLM and BERT embeddings are similar. Given SentenceBert’s widespread acceptance for clustering tasks and computational efficiency, we chose it for our SMART filtering methodology.
5.3 Quality of Filtering
To ensure that our method effectively filters out unwanted instances, we manually looked at 100 removed examples from the Easy and Similar examples category. For Similar examples, we found that 93 were genuinely similar, 5% were borderline similar and 2% were not similar. For Filtering Easy Examples, 97 examples could be easily answered by the authors or found via a simple web search. This analysis shows the effectiveness of our approach as most of the filtered out examples are unwanted for comparing model performance.
6 Discussion
Dataset Quality.
Our approach removes the easy, data contaminated, and similar examples, thus improving the quality of the resulting dataset. Our methodology can be applied at various stages, including before or after dataset release. SMART filtering is designed to be iterative and scalable; we intend to periodically reapply SMART-Filtering to existing benchmarks to identify and filter out examples that become too easy or contaminated for newer models. We suspect that, as models become more capable, what constitutes an easy example could change.
Computational Efficiency.
Recent studies have explored testing models on dataset subsets while maintaining high correlation with original dataset performance Varshney et al. (2022). Our approach achieves substantial dataset size reductions, up to 68.9% for ARC, while preserving similar model rankings as shown in Tables 1 and 3. As dataset size correlates with computation time and evaluation costs, the 68.9% decrease in size should be reflected in similar decreases in time and evaluation cost. Our methodology thus offers significant time and cost savings for future benchmark evaluations, aligning with the growing need for efficient model assessment techniques.
Improved Correlation with ChatBot Arena.
We find that SMART filtering can lead to better correlations with Elo scores from ChatBot Arena Chiang et al. (2024). This suggests that model scores on SMART filtering subsets are more representative of real-world model performance. This finding highlights the potential for developing effective evaluation datasets that mimic real-world usage of the model without incurring substantial time and cost to capture human preferences.
Datasets are Not Yet Saturated.
We found that model accuracies for all tested models drop significantly on datasets after SMART filtering, shown in Table 2. This indicates these datasets may not be ready for retirement yet, contrary to what leaderboards might suggest, and there remains room for improvement. Relatedly, iteratively applying SMART filtering over time can enable more dynamic evaluations (Dinan et al., 2019; Nie et al., 2020; Gehrmann et al., 2021; Kiela et al., 2021; Potts et al., 2021; Gehrmann et al., 2022; Margatina et al., 2023; Park et al., 2023; Graciotti et al., 2024), decreasing the impact of saturation by wringing more utility out of existing benchmarks.
7 Conclusion
In this work, we proposed SMART filtering, a methodology for identifying a challenging and high-quality subset of any benchmark, existing or new, that can better capture the capabilities of models and rank them. To achieve this we remove easy, data contaminated, and similar examples from a dataset. SMART filtering achieved significant increases in computational efficiency and better correlation with human preference data than the original datasets. We anticipate our approach will be useful for improving current benchmarking practices as well as for dataset creators to find high-quality subsets before dataset release in the future.
8 Limitations
In this work we present a methodology that can be applied to any NLP task, However, the method we have used for identifying data contaminated examples, may not be directly applicable to non-question answering datasets. Additionally, we tried to identify and remove incorrect ground truth annotations from the dataset (see Appendix for details). However, our initial attempt did not yield satisfactory results, highlighting the need for more effective strategies to address this challenge. Consequently, if a dataset contains a significant number of annotation errors, the proportion of such examples may increase in the resulting SMART filtering datasets.
Appendix A Appendix
A.1 Finding Threshold for Similar Examples
To identify similar examples, we analyze cosine distances between SentenceBert embeddings of the examples and identify the first local maxima from the distribution (see Figure 3). Notably, we found that considering only the $100$ nearest neighbors for each example is sufficient to determine the threshold accurately (Figure 5). Our analysis, detailed in the Appendix, shows that threshold values obtained using different $k$ values for $k$-nearest neighbors converge after 100 neighbors, aligning closely with results from the entire dataset. For computational efficiency, we therefore use $100$ nearest neighbors to establish the threshold value.
A.2 Attempt to Identify Wrong Ground Truth
Benchmark datasets often contain wrong ground truths. In an attempt to identify such examples, we tried an algorithmic framework. We hypothesized that if all top-performing models icorrectly predicted an example with very high confidence (>0.8), then that example would likely be wrongly annotated. However, upon applying this method to the MMLU dataset, our hypothesis didn’t work well. Approximately 1.4% of MMLU examples were filtering using this criteria, but manual inspection revealed that nearly half of these flagged examples actually had correct ground truth annotations, despite unanimous high-confidence wrong predictions by the models. Thus we didn’t include this step in SMART filtering as it would require human involvement and may not provide sufficient efficiency gains to justify the additional resource expenditure.
Figure 5: Distribution of cosine distances between SentenceBert embeddings for different values of k nearest neighbors for each example in the MMLU dataset. We find that threshold values obtained using different $k$ values for $k$-nearest neighbors converge after 100 neighbors.
A.3 Exact Match Duplicates
Dataset creators often preprocess their datasets to remove duplicate examples before releasing them. Nevertheless, exact-match duplicates can still be present. Our analysis reveals that 1.2% of the MMLU dataset consists of identical questions, while ARC contains 0.2% duplicates. In contrast, CommonsenseQA does not have any exact duplicates, with a rate of 0%.
A.4 Category Wise Results
MMLU has 57 different categories and we found that different categories in MMLU are affected differently by SMART filtering. The percentage of examples removed by SMART filtering for each category is given in Table 9.
A.5 LLM Based Embedding for Similar Examples
In addition to using LLM2Vec embeddings, we also used the last token’s LLM embeddings as representation of entire sentences Neelakantan et al. (2022); Ma et al. (2024) The results as shown in Table 8, shows that SentenceBert embeddings are very similar to LLM based embeddings for identifying and removing similar examples.
Embedding Pair Percentage Overlap
SBert and Llama-3-70B 94.7
SBert and Qwen2-72B 95.2
SBert and Llama-3-8B 95.8
Table 8: Percentage of examples with extremely high semantic similarity, based on a distance in the embedding space smaller than a $\delta$. We define the process for computing $\delta$ in Section 3.4.
A.6 Computation Resources
For all experiments for this work, we utilized A100 80GB GPUs. Depending on the model evaluated, we used 1,2,3 or 4 GPUs for inferences. These GPUs were assembled in a cluster of 8 GPUs in a node. The cumulative computing time required to evaluate all the language models and complete the experiments amounted to approximately 2000 GPU hours. We also used LLMs for coding assistance for building our codebase.
A.7 Detailed Model Results
Table 10, 11, and 12 shows accuracy for all tested models on ARC, MMLU, and CommonsenseQA dataset and their SMART filtered versions.
Category Name Original # of examples SMART filtered examples Percentage of examples removed
Moral Scenarios 895 0 100.00%
High School Government And Politics 193 52 73.06%
Us Foreign Policy 100 31 69.00%
High School Psychology 545 182 66.61%
Miscellaneous 783 274 65.01%
Marketing 234 86 63.25%
Sociology 201 76 62.19%
International Law 121 46 61.98%
High School World History 237 91 61.60%
World Religions 171 70 59.06%
High School Us History 204 85 58.33%
High School Geography 198 83 58.08%
Medical Genetics 100 43 57.00%
High School Biology 310 134 56.77%
High School European History 165 72 56.36%
Logical Fallacies 163 72 55.83%
Management 103 46 55.34%
College Medicine 173 80 53.76%
Prehistory 324 155 52.16%
College Biology 144 70 51.39%
High School Microeconomics 238 121 49.16%
Clinical Knowledge 265 136 48.68%
Jurisprudence 108 56 48.15%
Astronomy 152 79 48.03%
Security Studies 245 129 47.35%
Computer Security 100 54 46.00%
Human Sexuality 131 72 45.04%
High School Computer Science 100 55 45.00%
Professional Psychology 612 337 44.93%
Nutrition 306 171 44.12%
High School Macroeconomics 390 228 41.54%
Public Relations 110 67 39.09%
Human Aging 223 138 38.12%
Business Ethics 100 62 38.00%
Professional Medicine 272 173 36.40%
Philosophy 311 199 36.01%
Moral Disputes 346 223 35.55%
Anatomy 135 89 34.07%
Conceptual Physics 235 163 30.64%
College Physics 102 75 26.47%
Electrical Engineering 145 107 26.21%
Virology 166 127 23.49%
High School Chemistry 203 158 22.17%
College Chemistry 100 79 21.00%
High School Statistics 216 173 19.91%
Machine Learning 112 90 19.64%
Econometrics 114 93 18.42%
College Computer Science 100 83 17.00%
Professional Accounting 282 244 13.48%
Elementary Mathematics 378 330 12.70%
Formal Logic 126 110 12.70%
Professional Law 1534 1340 12.65%
High School Physics 151 143 5.30%
High School Mathematics 270 259 4.07%
Abstract Algebra 100 96 4.00%
Global Facts 100 96 4.00%
College Mathematics 100 97 3.00%
Table 9: Results of each of the 57 categories of SMART filtering on MMLU dataset. Different categories are affected differently.
Model Name Accuracy on ARC Accuracy on ARC-SMART
Qwen2-72B-Instruct 0.939 0.83
Meta-Llama-3.1-70B-Instruct 0.934 0.819
Meta-Llama-3-70B-Instruct 0.933 0.819
gemma-2-27b-it 0.925 0.788
Phi-3-medium-4k-instruct 0.923 0.781
Phi-3.5-MoE-instruct 0.92 0.785
Mixtral-8x22B-Instruct-v0.1 0.916 0.762
gemma-2-9b-it 0.91 0.757
Yi-34B-Chat 0.909 0.745
Qwen1.5-32B-Chat 0.907 0.752
dbrx-instruct 0.905 0.732
Yi-34B 0.902 0.724
Yi-1.5-9B-Chat 0.894 0.728
Meta-Llama-3-8B-Instruct 0.883 0.721
Qwen2-7B-Instruct 0.882 0.697
Mixtral-8x7B-Instruct-v0.1 0.876 0.681
Mixtral-8x7B-v0.1 0.876 0.688
internlm2_5-20b-chat 0.875 0.675
internlm2_5-7b-chat 0.861 0.647
Llama-2-70b-hf 0.855 0.644
gemma-7b 0.837 0.611
Mistral-7B-v0.3 0.802 0.563
Mistral-7B-Instruct-v0.2 0.793 0.581
Qwen-7B-Chat 0.764 0.518
gemma-7b-it 0.738 0.531
Qwen-7B 0.733 0.476
falcon-40b 0.722 0.51
falcon-40b-instruct 0.722 0.507
OLMo-1.7-7B-hf 0.639 0.435
Table 10: Accuracy of different models on ARC and ARC-SMART filtering datasets. Almost all the models have similar ranking among the two versions of ARC.
Model Name Accuracy on MMLU Accuracy on MMLU-SMART
Qwen2-72B-Instruct 0.841 0.743
Meta-Llama-3.1-70B-Instruct 0.823 0.714
Meta-Llama-3-70B-Instruct 0.803 0.692
Phi-3.5-MoE-instruct 0.787 0.67
Phi-3-medium-4k-instruct 0.781 0.656
Mixtral-8x22B-Instruct-v0.1 0.779 0.653
Yi-1.5-34B-Chat 0.764 0.634
gemma-2-27b-it 0.762 0.639
Yi-34B 0.758 0.624
Qwen1.5-32B-Chat 0.754 0.615
Yi-34B-Chat 0.747 0.603
dbrx-instruct 0.732 0.6
gemma-2-9b-it 0.725 0.588
internlm2_5-7b-chat 0.711 0.568
Mixtral-8x7B-Instruct-v0.1 0.706 0.565
Mixtral-8x7B-v0.1 0.704 0.568
Qwen2-7B-Instruct 0.702 0.564
internlm2_5-20b-chat 0.701 0.567
Yi-1.5-9B-Chat 0.699 0.556
Llama-2-70b-hf 0.69 0.544
Meta-Llama-3-8B-Instruct 0.664 0.505
gemma-7b 0.644 0.492
Mistral-7B-v0.3 0.626 0.468
Mistral-7B-Instruct-v0.2 0.593 0.441
Qwen-7B 0.574 0.426
Qwen-7B-Chat 0.563 0.415
falcon-40b 0.558 0.412
falcon-40b-instruct 0.547 0.402
OLMo-1.7-7B-hf 0.521 0.381
gemma-7b-it 0.517 0.389
Table 11: Model Accuracy Comparison on MMLU and MMLU-SMART filtering. Almost all the models have similar ranking among the two versions of MMLU.
Model Name Accuracy on CommonsenseQA Accuracy on CommonsenseQA-SMART
Qwen2-72B-Instruct 0.885 0.845
Yi-1.5-34B-Chat 0.833 0.776
Meta-Llama-3-70B-Instruct 0.833 0.771
Qwen1.5-32B-Chat 0.830 0.767
Meta-Llama-3.1-70B-Instruct 0.815 0.741
Phi-3.5-MoE-instruct 0.809 0.739
gemma-2-9b-it 0.808 0.733
Qwen2-7B-Instruct 0.804 0.724
Phi-3-medium-4k-instruct 0.797 0.722
Yi-34B 0.797 0.718
gemma-2-27b-it 0.794 0.719
Yi-34B-Chat 0.791 0.712
Yi-1.5-9B-Chat 0.790 0.718
internlm2_5-7b-chat 0.789 0.714
dbrx-instruct 0.786 0.704
internlm2_5-20b-chat 0.777 0.695
Meta-Llama-3-8B-Instruct 0.765 0.680
Mixtral-8x22B-Instruct-v0.1 0.759 0.672
OLMo-1.7-7B-hf 0.739 0.670
Mixtral-8x7B-Instruct-v0.1 0.706 0.600
gemma-7b-it 0.688 0.594
Qwen-7B 0.683 0.586
Mistral-7B-Instruct-v0.2 0.682 0.590
falcon-40b-instruct 0.678 0.579
Qwen-7B-Chat 0.652 0.557
gemma-7b 0.646 0.551
Mistral-7B-v0.3 0.603 0.499
Llama-2-70b-hf 0.583 0.465
Mixtral-8x7B-v0.1 0.581 0.468
falcon-40b 0.555 0.446
Table 12: Accuracy comparison of models on CommonsenseQA and CommonsenseQA-SMART filtering. Almost all the models have similar ranking among the two versions of CommonsenseQA.